If tanA + 1/tanA =2 then find value of tan^10A + 1/tan^10A

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Remember
X + 1/x is always gretaer than or equal to 2.
It is equal to 2 only when x = 1
Hence.
Tan A = 1 in this case
 
Now
tan ^10A + 1/tan^10 A = 1 + 1 = 2
 
Regards
Arun (askIITians forum expert)